# Image Generation Models: A Technical History

### 08 Mar 2026 v1

@@ 5.2 Transformer Autoregressive Image Generation
After the Transformer architecture [66] was introduced in 2017, it was shown that these architectures have strong ability to capture long term dependency as well as training parallelization. Although the transformer architecture was originally proposed for text and natural language applications, the architecture was eventually adopted for image applications. Parmar et al. [70] utilized the transformer architecture for auto-regressive image generation. The underlying principle is the same, given the pixels generated so far, what is the most likely next pixel.
@@
Inspired by autoregressive unsupervised learning for natural language data, Chen et al. [71] introduced iGPT. They train a transformer-style model to autoregressively predict pixels by treating the image as a 1-D token sequence, as shown in Fig. 40. Each pixel’s RGB is mapped to the nearest of 512 learned palette colors. The image is first downsampled to $48\times 48$ which results in a sequence of length $T=2304$ with $x_{t}\in\{0,\ldots,511\}$ for $t=1,\ldots,T$. Using the low-resolution and palette trick, they were able to decrease the context from $(224\times 224\times 3)$ to $64\times 64$ at most. The rest of the process is similar to how a transformer-based model is trained. The process includes token/position embeddings, normalization layers, transformer blocks, and output heads.
@@
Using a causal mask, they minimize the negative log-likelihood
$\mathcal{L}=-\sum_{t=1}^{T}\log p(x_{t}\mid x_{<t}).$ (72)
They pretrained autoregressive transformers on pixels and evaluated representations via linear probing/fine-tuning for classification. They normalized features across positions to get $f$, they then trained a linear classifier on the final feature. Using this, they were able to beat state-of-the-art unsupervised transfer methods at the time.
@@ 5.3 Scaling Up Text to Image Transformers
Although Chen et al. [71] did not illustrate how iGPT can be utilized for image generation, they provided a great foundation for the image generation research that followed. Ramesh et al. [72] introduced zero-shot text-to-image generation (DALL-E 1). Applying transformers at the pixel level is computationally intractable, especially for large images. Instead, they decided to represent an image as a composition of rich image constituents. Each of these constituents can be treated in the transformer model by its own token. Using this approach, the transformer model is able to capture the global correlation between different blocks within the image. For training, they utilize a two-stage approach that combines VQVAE [1] in stage 1 with a transformer in stage 2. Their approach relies on learning codebook vectors, $Z$, in stage 1 that act as a tokenizer for image partitions. In stage 1, they train a VQVAE [1] to compress the $256\times 256$ RGB image into a $32\times 32$ grid of image tokens. In stage 1, the model learns a codebook, an encoder, and a decoder. For a more detailed description of VQVAEs, please refer to the VAE section of this article. For capturing local dependencies they utilize CNN in the encoder and decoder parts of the model. In stage 2, they freeze the VQVAE and concatenate up to 256 text tokens (from image captions) with the $32\times 32=1024$ image tokens. Using this setup, they train an autoregressive transformer architecture to model the joint distribution between the image text description and the image itself. During generation (inference), they start with the text description. Using text tokens, they autoregressively generate $32\times 32$ image tokens. These tokens are then passed to the codebook and then the VAE decoder for the final generated image.
The transformer model used in stage 2 was a massive 12B-parameter model. In order to expedite training for this massive scale, they utilized mixed-precision coupled with distributed training.
Fig. 41 shows some of the sample images generated from text captions on the CUB dataset [73]. Note that these examples are zero-shot since the model was never trained on this dataset.
@@
In phase 2, the VQGAN is frozen and image $x$ is passed to the encoder and then vector quantizer to get code indices $s\in\{0,\ldots,|Z|-1\}^{h\times w}$, where $h,w$ are downsampling factors (e.g., $16\times 16$). Now a decoder-only autoregressive transformer is trained on these sequences $s$, where $N$ is the total number of patches (e.g. $16\times 16=256$):
$\log p(s)=\sum_{i=1}^{N}\log p(s_{i}\mid s_{<i}).$ (73)
@@
The same group of authors also introduced a follow up improvement and named it Muse [78]. They used a similar masked approach to MaskGIT [77], but made several architectural changes, as shown in Fig. 47. Instead of training a separate text encoder, they start from a pretrained language model that has good text understanding, T5 [79]. Moreover, they increased the number of VQGAN tokenizers from one to two: a base version to fill $16\times 16$ masked tokens and a high resolution model to map $16\times 16$ to $64\times 64$ tokens.
@@ 5.4 Conclusion
Autoregressive image generation predicts each pixel from the previously generated pixels. Unlike VAEs and GANs, these models are trained by maximum likelihood with a tractable objective. Moreover, they support conditioning naturally via the autoregressive factorization and masking. These architectures can easily incorporate pretrained text encoders for conditional generation, which makes them strong at contextual understanding. They come with downsides such as $O(n^{2})$ attention cost in token length. In the next section, we are going to review diffusion-based models and how they can generate even higher-quality images.
@@ 6 Diffusion Based Models
Unlike image generation models (VAEs, GANs, Normalizing Flows, Transformers) that we have investigated so far, diffusion models draw inspiration from physics. For example, when a sugar cube is dropped into a cup of tea, the sugar molecules gradually diffuse into the water, and this diffusion process can be described as a stochastic process. Diffusion models learn to reverse such a stochastic diffusion process where they transform noise into data. In the original image diffusion model proposed by Sohl-Dickstein et al. [82], a similar idea is used: starting from a clean image, a small amount of Gaussian noise is added at consecutive time steps until the image becomes nearly indistinguishable from isotropic Gaussian noise $\mathcal{N}(0,I)$. The goal of a diffusion model is to learn the reverse of this noising process, transforming pure noise back into the original image through a learned reverse diffusion procedure.
@@
Our goal is to train a model with parameters $\theta$ that learns the denoising transition process $p_{\theta}\bigl(x_{(t-1)}\mid x_{(t)}\bigr)$. By taking the log-likelihood of the generative model and applying Jensen’s inequality to the expectation over forward trajectories, we can obtain a lower bound $K$ on the log-likelihood:
@@
In other words, they add noise at step $t$ and try to predict the noise with an MSE loss. They also make significant architectural improvements. They use a UNet style CNN [85], coupled with downsampling and upsampling paths. Additionally, they incorporate self-attention at the lower feature-map resolution $(16\times 16)$. Box 8 shows the overall training and inference algorithm for a diffusion model trained via DDPM.
@@ 6.1 Further Improvements to Diffusion Models
where $z\sim\mathcal{N}(0,I)$.
$\sigma_{t}$ controls the randomness at each step. A fully stochastic version corresponds to the classic DDPM, and a less stochastic version is DDIM. Two consequences of these changes are that $\sigma_{t}$ acts as a control knob for sampling stochasticity. Because the model is trained to denoise any $t$, DDIM can also sample using a subset of timesteps (skipping steps) to accelerate generation. During training, the model only sees the forward noising process
@@
Despite strong sample quality, many diffusion models lagged behind GANs in terms of log-likelihood. Log-likelihood captures something more fundamental: how well the model matches the data distribution. Nichol and Dhariwal [87] add a variational lower-bound term to the loss and give the model more freedom (via learned variance) in order to achieve a tighter bound. Their model outputs an extra variance head $v=v_{\theta}(x_{t},t)$, and they turn this into variances as follows:
@@
where $L_{\text{vlb}}$ is the variational lower-bound loss. In practice, they set $\lambda$ to a small value so that $L_{\text{vlb}}$ guides the variance, but $L_{\text{simple}}$ is still the main source of influence on the output image. As a result of this change, they directly optimize the log-likelihood bound (VLB). Additionally, they introduce a cosine scheduling scheme for $\beta_{t}$ instead of linear scheduling. With this change, the noise addition is spread more smoothly across timesteps and this leads to more robust sampling with fewer time steps. Table VI shows the comparison between negative log-likelihood in their work and several other generative models on ImageNet and CIFAR datasets. Simultaneously, Song et. al. [88] introduced a score based continues approach to generative modeling that relied on Stochastic Differential Equations (SDEs). They utilize SDE to transform the data distribution into a known prior distribution (noise) and a corresponding SDE for the reverse process. They were able to achieve a competitive log-likelihood of 2.99 on CIFAR as presented in the table below.
@@
Salimans and Ho [91] introduce the idea of distillation for diffusion models with the goal of dramatically reducing the number of denoising steps. They start from a pretrained teacher diffusion model that must be sampled for $2N$ steps to go from pure noise to an image. The goal is to train a student model with the exact same architecture that can produce the same image as the teacher in only $N$ steps; i.e., each student step imitates two teacher steps.
@@
Yin et al. [92] achieve a $100\times$ speed-up in generation by introducing Distribution Matching Distillation (DMD), which approximately minimizes a KL divergence between two diffusion models, one tracking the generator’s ”fake” distribution and a fixed teacher model for the ”real” distribution. Later, Song et al. [93] introduced consistency models for fast one step generation while still allowing multi-step generation for more quality.
As shown in Fig. 53, they train a function $f_{\theta}$ such that any noisy sample $x_{t}$ on the same trajectory maps back to the origin, i.e., $f_{\theta}(x_{t},t)\approx x_{0}$. This enables them to do one-step generation by sampling $x_{T}\sim\mathcal{N}(0,I)$ and outputting $\hat{x}_{0}=f_{\theta}(x_{T},T)$, where $\hat{x}_{0}$ is treated as the generated image.
@@ 6.2 Conditional Diffusion Models
Dhariwal and Nichol [95] make further improvements to DDPM models by updating the model architecture and classifier guidance for conditional image generation. For the model, they perform a large ablation study and come up with the Ablated Diffusion Model (ADM). They modify the model architecture in the following ways: increase the model depth versus width, increase the number of attention heads, use attention heads at $32\times 32$, $16\times 16$, and $8\times 8$ resolutions rather than only at $16\times 16$, and make a few more architectural changes to the residual blocks. For conditional image generation, they exploit a classifier $p_{\phi}(y\mid x_{t})$ to further improve diffusion generation. During sampling, they modify the reverse diffusion step to shift the mean along the classifier gradient $\nabla_{x_{t}}\log p_{\phi}(y\mid x_{t})$.
To condition on label $y$, they want to sample from
@@
$\tilde{\epsilon}_{\theta}(z_{t},c)=(1+w)\,\epsilon_{\theta}(z_{t},c)-w\,\epsilon_{\theta}(z_{t},\varnothing)$ (92)
@@ 6.3 Diffusion Models Transition into Latent Space
So far we talked about the diffusion models that operate directly on the pixel space, and this can make them costly and slow. Rombach et al. [101] instead apply diffusion models in the latent space. Like many successful image generation models, their approach consists of two stages. In stage 1, they train a perceptual autoencoder that learns to reconstruct the input image. The encoder down-samples by a factor $f$ (e.g., $4$ or $8$) to generate a latent $z$:
@@
In stage 2, they train a diffusion model in the latent space (LDM). On the latent domain $z$, they train a classic $\epsilon$-prediction diffusion model. They also add more flexibility to diffusion models by augmenting the UNet backbone with a cross-attention mechanism for conditional image generation applications. They preprocess the condition(s) $y$ from various modalities via a domain-specific encoder $T_{\theta}$ that projects $y$ into an intermediate representation $\tau_{\theta}(y)\in\mathbb{R}^{M\times d_{c}}$. This is mapped to intermediate layers of the UNet via cross-attention. The conditional latent diffusion loss becomes
@@
Figure 57: Latent diffusion model maps the input condition to the intermediate layers of UNet via cross attention and generates output images based on input condition. Here the input is conditioned on removing the highlighted parts of the image. Image source: [101].
@@
$\tilde{y}=\sqrt{\gamma}\,y+\sqrt{1-\gamma}\,\epsilon,\qquad\epsilon\sim\mathcal{N}(0,I).$ (96)
@@ 6.4 Scaling up Diffusion Models
Ramesh et al. [107] introduced hierarchical text-conditional image generation with CLIP latents, also known as DALL-E 2. They leverage a large pretrained CLIP model in order to guide the generation process in diffusion models, trained on roughly 650M image and caption pairs. They then freeze the CLIP model and train two subsequent models, $\theta$ and $\phi$. Given an input image $x$, the CLIP model produces $z_{\text{img}}$ and $z_{\text{txt}}$ for the corresponding image–caption pairs. The two components of the model consist of a prior $p_{\phi}(z_{\text{img}}\mid y)$ and a decoder $p_{\theta}(x\mid z_{\text{img}},y)$ that produces images $x$ given the image embedding $z_{\text{img}}$ and optionally caption $y$:
@@
Simplified Training Algorithm for DALL-E 2 (un-CLIP) 1. Train a CLIP model on image–caption pairs $(x,y)$. 2. Freeze the CLIP model and compute image and text embeddings via $z_{\mathrm{img}}=E_{\mathrm{CLIP,img}}(x),\qquad z_{\mathrm{txt}}=E_{\mathrm{CLIP,txt}}(y).$ 3. Train the diffusion decoder $p_{\theta}(x\mid z_{\mathrm{img}},y)$ for NN iterations. For each mini-batch (size $m$): Compute CLIP embeddings $z_{\mathrm{img}}^{(i)},z_{\mathrm{txt}}^{(i)}$ from the input images $x^{(i)}$. Sample diffusion time steps and noise: $t\sim\mathrm{Uniform}\{1,\dots,T\},\qquad\epsilon^{(i)}\sim\mathcal{N}(0,I).$ Perform the forward noising process to obtain $x_{t}^{(i)}$. Do a forward pass through the decoder UNet: $\hat{\epsilon}^{(i)}=\epsilon_{\theta}\bigl(x_{t}^{(i)},t,z_{\mathrm{img}}^{(i)},z_{\mathrm{txt}}^{(i)}\bigr)$ Compute the noise-prediction loss and update weights using $L_{\mathrm{decoder}}=\frac{1}{m}\sum_{i}\left\lVert\epsilon^{(i)}-\hat{\epsilon}^{(i)}\right\rVert_{2}^{2}$ 4. Train the diffusion prior $p_{\phi}(z_{\mathrm{img}}\mid y)$ for $N^{\prime}$ iterations. For each mini-batch (size $m$): Compute CLIP embeddings $z_{\mathrm{img}}^{(i)},z_{\mathrm{txt}}^{(i)}$. Sample diffusion time steps and noise: $t\sim\mathrm{Uniform}\{1,\dots,T\},\ \epsilon^{(i)}\sim\mathcal{N}(0,I)$ Generate noisy CLIP image embeddings via the forward noising process $z_{\mathrm{img},t}^{(i)}=\text{forward\_noising}\bigl(z_{\mathrm{img}}^{(i)},t,\epsilon^{(i)}\bigr)$ Run the reverse diffusion process with the prior model $\hat{z}_{\mathrm{img}}^{(i)}=f_{\phi}\bigl(z_{\mathrm{img},t}^{(i)},t;\epsilon^{(i)}\bigr)$ Compute the denoising loss and update weights by $\nabla_{\phi}L_{\mathrm{prior}}$: $L_{\mathrm{prior}}=\frac{1}{m}\sum_{i}\left\lVert\hat{z}_{\mathrm{img}}^{(i)}-z_{\mathrm{img}}^{(i)}\right\rVert_{2}^{2}$
@@
Later, Saharia et al. [108] introduced Imagen, where the model is architecturally simpler than DALL-E 2 but uses a very large pretrained T5 [79] model (T5-XXL) as the text encoder. They use three diffusion models as depicted in Fig. 63: one base diffusion model to generate a $64\times 64$ image, a second model to upsample images to $256\times 256$, and a final diffusion model to upsample to $1024\times 1024$. The base diffusion model consists of an efficient UNet–style noise predictor
@@
where $w(t)$ is a cosine-based weight schedule. At the end, they use the two additional diffusion models to further upsample the generated $64\times 64$ image to $1024\times 1024$.
@@
Table VIII compares zero-shot FID scores between Imagen and other state of the art models. As you can see, Imagen significantly improves the FID scores on the MS-COCO dataset compared to other models. Fig. 64 compares a few sample images generated by Imagen and DALL-E 2.
@@
At approximately the same time, Podell and coauthors [110] introduced Stable Diffusion XL (SDXL) for high-resolution image synthesis. Compared to Stable Diffusion v1 or v2, they utilize a UNet backbone that is roughly three times larger. They also use richer text conditioning by concatenating the outputs of two encoders: CLIP ViT-L [99] and OpenCLIP ViT-bigG. In addition, they include a refinement model at the end of the pipeline that is responsible for generating the final high-resolution images. Fig. 66 shows the high-level diagram of their approach.
@@
Table X shows a comparison between SDXL and older Stable Diffusion (SD) models v1.x and v2.x. Fig. 67 shows sample images generated by SDXL and compares them with Stable Diffusion (SD) v1.5 and v2.1.

### 29 Mar 2026 v2

@@ 5.2 Transformer Autoregressive Image Generation
After the Transformer architecture [66] was introduced in 2017, it was shown that these architectures have strong ability to capture long term dependency as well as training parallelization. Although the transformer architecture was originally proposed for text and natural language applications, the architecture was eventually adopted for image applications. Parmar et al. [70] utilized the transformer architecture for auto-regressive image generation, Image Transformer. The underlying principle is the same, given the pixels generated so far, what is the most likely next pixel.
@@
Inspired by autoregressive unsupervised learning for natural language data, Chen et al. [71] introduced iGPT. They train a transformer-style model to autoregressively predict pixels by treating the image as a 1-D token sequence, as shown in Fig. 40. Each pixel’s RGB is mapped to the nearest of 512 learned palette colors. The image is first downsampled to $48\times 48$ which results in a sequence of length $T=2304$ with $x_{t}\in\{0,\ldots,511\}$ for $t=1,\ldots,T$. Using the low-resolution and palette trick, they were able to decrease the context from $(224\times 224\times 3)$ to $64\times 64$ at most. The rest of the process is similar to how a transformer-based model is trained. The process includes token/position embeddings, normalization layers, transformer blocks, and output heads. Using a causal mask, they minimize the negative log-likelihood:
$\mathcal{L}=-\sum_{t=1}^{T}\log p(x_{t}\mid x_{<t}).$ (72)
They pretrained an autoregressive transformer on pixels and evaluated representations via linear probing/fine-tuning for classification. They normalized features across positions and then trained a linear classifier on the final feature. Using this, they were able to beat state-of-the-art unsupervised transfer methods at the time.
@@ 5.3 Scaling Up Text to Image Transformers
Although Chen et al. [71] did not illustrate how iGPT can be utilized for image generation, they provided a great foundation for the image generation research that followed. Ramesh et al. [72] introduced zero-shot text-to-image generation (DALL-E 1). Applying transformers at the pixel level is computationally intractable, especially for large images. Instead, they decided to represent an image as a composition of rich image constituents. Each of these constituents can be treated in the transformer model by its own token. Using this approach, the transformer model is able to capture the global correlation between different blocks within the image. For training, they utilize a two-stage approach that combines VQ-VAE [1] in stage 1 with a transformer in stage 2. Their approach relies on learning codebook vectors, $Z$, in stage 1 that act as a tokenizer for image partitions. In stage 1, they train a VQ-VAE [1] to compress the $256\times 256$ RGB image into a $32\times 32$ grid of image tokens. In stage 1, the model learns a codebook, an encoder, and a decoder. For a more detailed description of VQ-VAEs, please refer to section 2.6.3 in the VAE section of this article. For capturing local dependencies they utilize CNN in the encoder and decoder parts of the model. In stage 2, they freeze the VQ-VAE and concatenate up to 256 text tokens (from image captions) with the $32\times 32=1024$ image tokens. Using this setup, they train an autoregressive transformer architecture to model the joint distribution between the image text description and the image itself. During generation (inference), they start with the text description. Using text tokens, they autoregressively generate $32\times 32$ image tokens. These tokens are then passed to the codebook and then the VAE decoder for the final generated image.
The transformer model used in stage 2 was a massive 12B-parameter model. In order to expedite training for this massive scale, they utilized mixed-precision coupled with distributed training. Fig. 41 shows some of the sample images generated from text captions on the CUB dataset [73]. Note that these examples are zero-shot since the model was never trained on this dataset.
In phase 2, the VQGAN is frozen and image $x$ is passed to the encoder and then vector quantizer to get code indices $s\in\{0,\ldots,|Z|-1\}^{h\times w}$, where $h,w$ are downsampling factors (e.g., $16\times 16$). Now a decoder-only autoregressive transformer is trained on these sequences $s$, where NN is the total number of patches (e.g. $16\times 16=256$):
$\log p(s)=\sum_{i=1}^{N}\log p(s_{i}\mid s_{<i}).$ (73)
@@
The same group of authors also introduced a follow up improvement and named it Muse [78]. They used a similar masked approach to MaskGIT [77], but made several architectural changes, as shown in Fig. 47. Instead of training a separate text encoder, they start from a pretrained language model that has good text understanding, T5 [79]. Moreover, they increased the number of VQ-GAN tokenizers from one to two: a base version to fill $16\times 16$ masked tokens and a high resolution model to map $16\times 16$ to $64\times 64$ tokens.
@@ 5.4 Conclusion
Autoregressive image generation predicts each pixel from the previously generated pixels. Unlike VAEs and GANs, these models are trained by maximum likelihood with a tractable objective. Moreover, they support conditioning naturally via the autoregressive factorization and masking. These architectures can easily incorporate pretrained text encoders for conditional generation, which makes them strong at contextual understanding. Despite all of these benefits, they come with downsides such as $O(n^{2})$ attention cost in token length. In the next section, we are going to review diffusion-based models and how they can generate even higher-quality images.
@@ 6 Diffusion Based Models
Unlike image generation models that we have investigated so far (VAEs, GANs, Normalizing Flows, Transformers), diffusion models draw inspiration from physics. For example, when a sugar cube is dropped into a cup of tea, the sugar molecules gradually diffuse into the water, and this diffusion process can be described as a stochastic process. Diffusion models learn to reverse such a stochastic diffusion process where they transform noise into data. In the original image diffusion model proposed by Sohl-Dickstein et al. [82], a similar idea is used: starting from a clean image, a small amount of Gaussian noise is added at consecutive time steps until the image becomes nearly indistinguishable from isotropic Gaussian noise $\mathcal{N}(0,I)$. The goal of a diffusion model is to learn the reverse of this noising process, transforming pure noise back into the original image through a learned reverse diffusion procedure.
@@
Our goal is to train a model (with parameters $\theta$) that learns the denoising transition process $p_{\theta}\bigl(x_{(t-1)}\mid x_{(t)}\bigr)$. By taking the log-likelihood of the generative model and applying Jensen’s inequality to the expectation over trajectories, we can obtain a lower bound $K$ on the log-likelihood:
@@
They also make significant architectural improvements. They use a UNet style CNN [85], coupled with downsampling and upsampling paths. Additionally, they incorporate self-attention at the lower feature-map resolution $(16\times 16)$. Box 8 shows the overall training and inference algorithm for a diffusion model trained via DDPM.
@@ 6.1 Further Improvements to Diffusion Models
where $z\sim\mathcal{N}(0,I)$. $\sigma_{t}$ controls the randomness at each step. A fully stochastic version corresponds to the classic DDPM, and a less stochastic version is DDIM. Two consequences of these changes are that $\sigma_{t}$ acts as a control knob for sampling stochasticity. Because the model is trained to denoise any $t$, DDIM can also sample using a subset of timesteps (skipping steps) to accelerate generation. During training, the model only sees the forward noising process
@@
Despite strong sample quality, many diffusion models lagged behind GANs in terms of log-likelihood. Log-likelihood captures something more fundamental: how well the model matches the data distribution. Nichol and Dhariwal [87] added a variational lower-bound term to the loss and gave the model more freedom (via learned variance) in order to achieve a tighter bound. Their model outputs an extra variance head $v=v_{\theta}(x_{t},t)$, and they turn this into variances as follows:
@@
where $L_{\text{vlb}}$ is the variational lower-bound loss. In practice, they set $\lambda$ to a small value so that $L_{\text{vlb}}$ guides the variance, but $L_{\text{simple}}$ is still the main source of influence on the output image. As a result of this change, they directly optimize the log-likelihood bound (VLB). Additionally, they introduce a cosine scheduling scheme for $\beta_{t}$ instead of linear scheduling. With this change, the noise addition is spread more smoothly across timesteps and this leads to more robust sampling with fewer time steps. Table VI shows the comparison between negative log-likelihood in their work and several other generative models on ImageNet and CIFAR datasets. Simultaneously, Song et al. [88] introduced a score based continues approach to generative modeling that relied on Stochastic Differential Equations (SDEs). They utilize SDE to transform the data distribution into a known prior distribution (noise) and a corresponding SDE for the reverse process. They were able to achieve a competitive log-likelihood of 2.99 on CIFAR as presented in the table below.
@@
Salimans and Ho [91] introduced the idea of distillation for diffusion models with the goal of dramatically reducing the number of denoising steps. They started from a pretrained teacher diffusion model that must be sampled for $2N$ steps to go from pure noise to an image. The goal is to train a student model with the exact same architecture that can produce the same image as the teacher in only NN steps; i.e., each student step imitates two teacher steps.
@@
Yin et al. [92] achieve a $100\times$ speed-up in generation by introducing Distribution Matching Distillation (DMD), which approximately minimizes a KL divergence between two diffusion models, one tracking the generator’s ”fake” distribution and a fixed teacher model for the ”real” distribution. Later, Song et al. [93] introduced consistency models for fast one step generation while still allowing multi-step generation for more quality. As shown in Fig. 53, they train a function $f_{\theta}$ such that any noisy sample $x_{t}$ on the same trajectory maps back to the origin, i.e., $f_{\theta}(x_{t},t)\approx x_{0}$. This enables them to do one-step generation by sampling $x_{T}\sim\mathcal{N}(0,I)$ and outputting $\hat{x}_{0}=f_{\theta}(x_{T},T)$, where $\hat{x}_{0}$ is treated as the generated image.
@@ 6.2 Conditional Diffusion Models
Dhariwal and Nichol [95] make further improvements to DDPM models by updating the model architecture and classifier guidance for conditional image generation. For the model, they perform a large ablation study and come up with the Ablated Diffusion Model (ADM). They modify the model architecture in the following ways: increase the model depth versus width, increase the number of attention heads, use attention heads at $32\times 32$, $16\times 16$, and $8\times 8$ resolutions rather than only at $16\times 16$, and make a few more architectural changes to the residual blocks. For conditional image generation, they exploit a separate classifier model $p_{\phi}(y\mid x_{t})$ to further improve diffusion generation. During sampling, they modify the reverse diffusion step to shift the mean along the classifier gradient $\nabla_{x_{t}}\log p_{\phi}(y\mid x_{t})$. To condition on label $y$, they want to sample from
@@
$\tilde{\epsilon}_{\theta}(x_{t},c)=(1+w)\,\epsilon_{\theta}(x_{t},c)-w\,\epsilon_{\theta}(x_{t},\varnothing)$ (92)
@@ 6.3 Diffusion Models Transition into Latent Space
So far we talked about the diffusion models that operate directly on the pixel space, and this can make them costly and slow. Rombach et al. [101] instead decided to apply diffusion models in the latent space. Like many successful image generation models, their approach consists of two stages. In stage 1, they train a perceptual autoencoder that learns to reconstruct the input image. The encoder down-samples by a factor $f$ (e.g., 44 or 88) to generate a latent $z$:
@@
In stage 2, they train a diffusion model in the latent space, a latent diffusion model (LDM). On the latent domain $z$, they train a classic $\epsilon$-prediction diffusion model. They also add more flexibility to diffusion models by augmenting the UNet backbone with a cross-attention mechanism for conditional image generation applications. They preprocess the condition(s) $y$ from various modalities via a domain-specific encoder $T_{\theta}$ that projects $y$ into an intermediate representation $\tau_{\theta}(y)\in\mathbb{R}^{M\times d_{c}}$. This is mapped to intermediate layers of the UNet via cross-attention. The conditional latent diffusion loss becomes
@@
Figure 57: Latent diffusion model (LDM) maps the input condition to the intermediate layers of UNet via cross attention and generates output images based on input condition. Here the input is conditioned on removing the highlighted parts of the image. Image source: [101].
@@
$\tilde{y}=\sqrt{\gamma}\,y+\sqrt{1-\gamma}\,\epsilon,\qquad\epsilon\sim\mathcal{N}(0,I)$ (96)
@@ 6.4 Scaling up Diffusion Models
Ramesh et al. [107] introduced hierarchical text-conditional image generation with CLIP latents, also known as DALL-E 2. They leverage a large pretrained CLIP model in order to guide the generation process in diffusion models, trained on roughly 650M image and caption pairs. They then freeze the CLIP model and train two subsequent models, $\theta$ and $\phi$. Given an input image $x$, the CLIP model produces $z_{\text{img}}$ and $z_{\text{txt}}$ for the corresponding image–caption pairs. The two components of the model consist of a prior $p_{\phi}(z_{\text{img}}\mid y)$ and a decoder $p_{\theta}(x\mid z_{\text{img}},y)$ that produces images $x$ given the embedding of the image $z_{\text{img}}$ and optionally caption $y$:
@@
Simplified Training Algorithm for DALL-E 2 (un-CLIP) 1. Train a CLIP model on image–caption pairs $(x,y)$. 2. Freeze the CLIP model and compute image and text embeddings via $z_{\mathrm{img}}=E_{\mathrm{CLIP,img}}(x),\qquad z_{\mathrm{txt}}=E_{\mathrm{CLIP,txt}}(y).$ 3. Train the diffusion decoder $p_{\theta}(x\mid z_{\mathrm{img}},y)$ for NN iterations. For each mini-batch (size $m$): Compute CLIP embeddings $z_{\mathrm{img}}^{(i)},z_{\mathrm{txt}}^{(i)}$ from the input images $x^{(i)}$. Sample diffusion time steps and noise: $t\sim\mathrm{Uniform}\{1,\dots,T\},\qquad\epsilon^{(i)}\sim\mathcal{N}(0,I).$ Perform the forward noising process to obtain $x_{t}^{(i)}$. Do a forward pass through the decoder UNet: $\hat{\epsilon}^{(i)}=\epsilon_{\theta}\bigl(x_{t}^{(i)},t,z_{\mathrm{img}}^{(i)},z_{\mathrm{txt}}^{(i)}\bigr)$ Compute the noise-prediction loss and update weights using $L_{\mathrm{decoder}}=\frac{1}{m}\sum_{i}\left\lVert\epsilon^{(i)}-\hat{\epsilon}^{(i)}\right\rVert_{2}^{2}$ 4. Train the diffusion prior $p_{\phi}(z_{\mathrm{img}}\mid y)$ for $N^{\prime}$ iterations. For each mini-batch (size $m$): Compute CLIP embeddings $z_{\mathrm{img}}^{(i)},z_{\mathrm{txt}}^{(i)}$. Sample diffusion time steps and noise: $t\sim\mathrm{Uniform}\{1,\dots,T\},\ \epsilon^{(i)}\sim\mathcal{N}(0,I)$ Generate noisy CLIP image embeddings via the forward noising process $z_{\mathrm{img},t}^{(i)}=\text{forward\_noising}\bigl(z_{\mathrm{img}}^{(i)},t,\epsilon^{(i)}\bigr)$ Run the reverse diffusion process with the prior model $\hat{z}_{\mathrm{img}}^{(i)}=f_{\phi}\bigl(z_{\mathrm{img},t}^{(i)},t;z_{txt}^{(i)}\bigr)$ Compute the denoising loss and update weights by $\nabla_{\phi}L_{\mathrm{prior}}$: $L_{\mathrm{prior}}=\frac{1}{m}\sum_{i}\left\lVert\hat{z}_{\mathrm{img}}^{(i)}-z_{\mathrm{img}}^{(i)}\right\rVert_{2}^{2}$
@@
Later, Saharia et al. [108] introduced Imagen, where the model is architecturally simpler than DALL-E 2 but uses a very large pretrained T5 [79] model (T5-XXL) as the text encoder. They use three diffusion models as depicted in Fig. 63: a base diffusion model to generate a $64\times 64$ image, a second model to upsample images to $256\times 256$, and a final diffusion model to upsample to $1024\times 1024$. The base diffusion model consists of an efficient UNet–style noise predictor
@@
where $w(t)$ is a cosine-based weight schedule. At the end, they use the two additional diffusion models to further upsample the generated $64\times 64$ image to $1024\times 1024$. Table VIII compares zero-shot FID scores between Imagen and other state of the art models. As you can see, Imagen significantly improves the FID scores on the MS-COCO dataset compared to other models. Fig. 64 compares a few sample images generated by Imagen and DALL-E 2.
@@
At approximately the same time, Podell and coauthors [110] introduced Stable Diffusion XL (SDXL) for high-resolution image synthesis. Compared to Stable Diffusion v1 or v2, they utilize a UNet backbone that is roughly three times larger. They also use richer text conditioning by concatenating the outputs of two encoders: CLIP ViT-L [99] and OpenCLIP ViT-bigG. In addition, they include a refinement model at the end of the pipeline that is responsible for generating the final high-resolution images. Fig. 66 shows the high-level diagram of their approach. Table X shows a comparison between SDXL and older Stable Diffusion (SD) models v1.x and v2.x. Fig. 67 shows sample images generated by SDXL and compares them with Stable Diffusion (SD) v1.5 and v2.1.